OpenStudy (anonymous):

16^32^64=2^2^X FIND X

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (watchmath):

\[16^{32^{64}}=16^{(2^5)^{64}}=16^{2^{320}}=(2^4)^{2^{320}}=2^{2^2\cdot 2^{320}}=2^{2^{322}}\] So \(x=322\)

OpenStudy (anonymous):

thank you :)